OpenStudy (anonymous):

yy''-y^'2+y'=0 general solution?

OpenStudy (jamesj):

Do you mean \[ yy'' - (y')^2 + y' = 0 \] ?

OpenStudy (anonymous):

yes

OpenStudy (mr.math):

wolframalpha suggests a nice substitution that makes the differential equation a separable one.

OpenStudy (jamesj):

Notice first that y y'' = y' (y' - 1) hence \[ y'' \left( \frac{1}{y'} - \frac{1}{y'-1} \right) = \frac{1}{y} \] Now integrate once, find a new equation involving only y' and y; and then solve again.

OpenStudy (jamesj):

I flipped the signs on the LHS.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

y'=P ,y''=PdP/dy,so P(ydP/dy-P+1)=O ..........

OpenStudy (jamesj):

Here's an elegant solution. Divide through by y^2 and note that \[ \left(\frac{y'}{y}\right)' = \frac{yy'' - (y')^2}{y^2} \ \ \hbox{ and } \ \ \left(\frac{1}{y}\right)'= -\frac{y'}{y^2} \]

OpenStudy (mr.math):

Very smart!